OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write the solution as an ordered pair.

OpenStudy (amistre64):

again ? lol

OpenStudy (amistre64):

m+5f = 25 m+ 10f = 40 -2m -10f = -50 m +10f = 40 --------------- -m = -10 m = 10 10 + 5f = 25 5f = 15 ; f=3